OpenStudy (reverdeparadis):

How do I graph this? -0.6x - 0.8y = 0.30

OpenStudy (solomonzelman):

\[-0.6x - 0.8y = 0.30~~~~~~~~~~~~~everything~~~\times 10,\]\[-6x - 8y =3\]\[-6x - 8y +6x=3+6x\]\[- 8y =3+6x\]\[- 8y =3+6x~~~~~~~~~~everything~~\div~(-8)\]\[y=-\frac{3}{4}x-\frac{3}{8}\]can you graph it now?